OpenStudy (anonymous):

The equation 1.6x - 2.2 = 1.4x can be transformed to form which of the following expressions? A. 3.0x = 2.2 B. 0.2x = -2.2 C. 3.0x = -2.2 D. 0.2x = 2.2

OpenStudy (anonymous):

@pooja195

OpenStudy (welshfella):

add 2.2 to both sides then add 1.6x to both sides

OpenStudy (welshfella):

** sorry the last line should be subtract 1.4x from both sides

OpenStudy (anonymous):

i dont understand @welshfella

OpenStudy (welshfella):

1.6x - 2.2 = 1.4x add 2.2 to both sides to get the number on right side 1.6x - 2.2 + 2,2 = 1.4x + 2.2 1.6x = 1.4x + 2.2 now get the 1.4x to left side by subtracting it from both sides 1.6x - 1.4x = 1.4x - 1.4x + 2.2 0.2x = 2.2

OpenStudy (anonymous):

so its d ? @welshfella